Suppose that in a certain precinct, 1150 people voted in the last presidential election. If the ratio of female voters to male voters was 2 to 3, how many females and how many males voted?
Answer by Dickiedannie(6)

I. ADD THE RATIO, 2+3=5
II.FEMALES WHO VOTED = (2/5)*1150= 460
III. MALES WHO VOTED = (3/5)*1150=690
THEREFORE THE FEMALES AND MALES WHO VOTED WERE 460 AND 690 RESPECTIVELY